OpenStudy (anonymous):

a-18/12=15/3 solve the proportion

OpenStudy (anonymous):

\[a=15/3+18/12\]

OpenStudy (anonymous):

Now you can do it yourself. 1. Find the LCM, 2. Then add the RHS thats all Does that help?

OpenStudy (anonymous):

a=6.5??

OpenStudy (anonymous):

Thats correct!

OpenStudy (anonymous):

ok cool

OpenStudy (gw2011):

Is this proportion meant to be: (a-18)/12=15/3 or a-(18/12)=15/3? Depending on how this proportion is set will make a difference in the answer.